# reference.bib

### v2

@@ VI CONCLUSION
Vision sensors are a critical component of autonomous vehicles, playing a key role in the decision-making process. As one of the fastest-growing fields in recent years, computer vision techniques are used to analyze data captured by vision sensors to obtain useful information such as traffic light detection, traffic sign recognition, drivable area detection, and 3D obstacle perception. With advancements in sensor technology, algorithmic capabilities, and computational power, the use of vision sensor data for autonomous vehicle perception is gaining increasing attention. Monocular-based 3D lane detection, for example, uses a single camera image to obtain the position of lane lines in the 3D physical world, incorporating depth information. Knowing the depth information of lane lines is crucial for safe and comfortable decision-making and planning in autonomous vehicles. While 3D lane information can be obtained using other sensors such as LiDAR, vision sensors play a crucial role in the field of autonomous driving due to their cost-effectiveness and rich structured color information.
@@
3D lane detection based on monocular image has been developed for many years in the field of autonomous driving. However, there is a lack of comprehensive and summarized analysis in the existing literature. In this survey, we reviewed existing lane detection methods, introduced existing three-dimensional lane detection datasets, and discussed the performance comparison of existing lane detection methods on public datasets. We also analyzed the challenges and limitations faced by current three-dimensional lane detection. Our main conclusion is that the research in the field of three-dimensional lane detection based on monocular images has not reached perfection, and current methods have many limitations, which we discussed in detail in the survey and provided relevant suggestions and outlooks. We covered baseline works dealing with deep learning models, their hierarchy for three-dimensional lane detection tasks, and the challenges associated with each model category. Furthermore, we delved into performance evaluation strategies, loss functions, and widely used datasets in the field of autonomous driving for three-dimensional lane detection models. We summarized the work by presenting open challenges and future research directions, along with stimulating baseline references from recent literature.
@@
Lastly, it is undeniable that experts from the intelligent transportation systems community are continuously striving to improve strategies for three-dimensional lane detection to effectively utilize data from vision sensors. Mainstream research is dedicated to enhancing model accuracy through the capabilities of neural networks or exploring novel neural network architectures. However, addressing other challenges is necessary to achieve reliable, trustworthy, and safe autonomous driving. From the perspective of three-dimensional lane detection, these challenges demand more robust models with the ability to anticipate lane occlusions, process coarse-structure information, and provide risk alerts. Additionally, current three-dimensional lane detection models primarily rely on supervised learning, which requires high-quality labeled data. However, annotating three-dimensional lane data is a laborious and time-consuming task. Exploring valuable and challenging approaches, such as self-supervised or weakly supervised learning, to achieve three-dimensional lane detection is an open opportunity for further development in this field. If these opportunities can be fully leveraged in a timely manner, it can drive research in intelligent transportation systems and elevate three-dimensional lane detection to a new level. This would enable unmanned vehicles to be deployed more effectively in real-world environments and support safer, more reliable, and more comfortable travel and logistics services.

### v3

@@ VI CONCLUSION
Vision sensors are essential in autonomous vehicles, as they significantly influence the decision-making process. As one of the fastest-evolving areas in recent years, computer vision techniques analyze data from these sensors to extract crucial information, including traffic light detection, traffic sign recognition, drivable area delineation, and 3D obstacle perception. With advancements in sensor technology, algorithmic sophistication, and computational power, leveraging vision sensor data for autonomous vehicle perception has garnered growing attention. Monocular-based 3D lane detection, for instance, uses a single camera image to identify lane line positions within the 3D environment, integrating valuable depth information. Depth knowledge of lane lines is essential for making safe, comfortable decisions and planning maneuvers. Although 3D lane information can also be derived from other sensors, such as LiDAR, vision sensors remain invaluable in AD for their cost-effectiveness and rich, structured color information.
@@
Monocular image-based 3D lane detection has been extensively studied in AD, however, comprehensive analysis and synthesis of this research are scarce in the current literature. In this survey, we systematically review existing lane detection methods, introduce available 3D lane detection datasets, and discuss the performance of these methods across public datasets. Additionally, we examine the challenges and limitations that current 3D lane detection approaches encounter. Our findings highlight that monocular image-based 3D lane detection remains an evolving field with notable limitations, which we discuss in detail along with corresponding recommendations and outlooks. We cover foundational work in deep learning models, outline their structural hierarchy for 3D lane detection tasks, and address the specific challenges encountered by each model category. Furthermore, we analyze performance evaluation strategies, loss functions, and widely used datasets in AD relevant to 3D lane detection models. Finally, we summarize open challenges and potential future research directions, providing reference to key baseline studies from recent literature to foster further progress in this area.
@@
Finally, it is clear that experts in the intelligent transportation systems field are continuously advancing strategies for 3D lane detection to fully harness data from vision sensors. Current research predominantly focuses on enhancing model accuracy via neural networks or by exploring innovative network architectures. However, addressing additional challenges remains essential to achieve reliable, trustworthy, and safe AD. Specifically, in 3D lane detection, these challenges demand the development of more robust models capable of anticipating lane occlusions, processing coarse-structured information, and issuing risk alerts. Moreover, existing 3D lane detection models largely depend on supervised learning, which necessitates high-quality labeled data—a labor-intensive and time-consuming requirement. Thus, exploring alternative methodologies such as self-supervised or weakly supervised learning presents a valuable and promising path for further development in this area. Effectively leveraging these opportunities could significantly propel research in intelligent transportation systems, advancing 3D lane detection to new heights. This progress would enable more effective deployment of autonomous vehicles in real-world settings, supporting safer, more reliable, and comfortable travel and logistics solutions.